### v3

# SEA: Low-Resource Safety Alignment for Multimodal Large Language Models via Synthetic Embeddings

Abstract
Multimodal Large Language Models (MLLMs) have serious security vulnerabilities. While safety alignment using multimodal datasets consisting of text and data of additional modalities can effectively enhance MLLM’s security, it is costly to construct these datasets. Existing low-resource security alignment methods, including textual alignment, have been found to struggle with the security risks posed by additional modalities. To address this, we propose Synthetic Embedding augmented safety Alignment (SEA), which optimizes embeddings of additional modality through gradient updates to expand textual datasets. This enables multimodal safety alignment training even when only textual data is available. Extensive experiments on image, video, and audio-based MLLMs demonstrate that SEA can synthesize a high-quality embedding on a single RTX3090 GPU within 24 seconds. SEA significantly improves the security of MLLMs when faced with threats from additional modalities. To assess the security risks introduced by video and audio, we also introduced a new benchmark called VA-SafetyBench. High attack success rates across multiple MLLMs validate its challenge. Our code and data will be available at https://github.com/ZeroNLP/SEA.
This paper contains harmful data and model-generated content that can be offensive in nature.
SEA: Low-Resource Safety Alignment for Multimodal Large Language Models via Synthetic Embeddings
footnotetext: *Corresponding author
1 Introduction
Multimodal Large Language Models (MLLMs) integrate additional modality encoders with large language models (LLMs), equipping them with the ability to comprehend and reason on multimodal data such as images Liu et al. (2024b, a); Chen et al. (2023), videos Wang et al. (2024b); Cheng et al. (2024), and audio Chu et al. (2024). Although MLLMs achieve advanced multimodal capability, they exhibit more serious security risks than LLMs. By injecting malicious information into non-textual inputs such as images Liu et al. (2024c); Li et al. (2024b) or audio Yang et al. (2024a), MLLMs can be easily induced to comply with users’ harmful instructions.
To address the aforementioned issues, current mitigation strategies, such as supervised fine-tuning (SFT) Zong et al. (2024) and reinforcement learning with human feedback (RLHF) Zhang et al. (2024) demonstrate effectiveness in enhancing the safety of MLLM. However, the construction of multimodal safety alignment datasets is costly. Unlike LLMs, high-quality safety alignment data for MLLMs requires a strong correlation between the three components: textual instructions, textual responses, and additional modalities, making the data collection process even more expensive. Moreover, due to differences in additional modalities, safety alignment data must be rebuilt whenever a new emerging modality (such as electroencephalogram signals Wang et al. (2024a)) is introduced for MLLM. This not only incurs additional costs but also causes the development of datasets to lag behind the advancements of the MLLMs themselves. Therefore, there is an urgent need for a resource-efficient and universally applicable safety alignment method to promote the development of safer MLLMs.
Recently, Chakraborty et al. (2024) revealed that textual alignment can significantly enhance the security of image-based MLLMs, providing a promising solution for low-resource safety alignment. However, further exploration by Hu et al. (2024) found that textual alignment is effective only when explicit harmful information appears in the text input, such as the instruction “how to use the product in the image to rob a bank” with an image input of a bomb. In contrast, models that have undergone multimodal alignment are generally effective across various scenarios, including samples that present harmful information solely through images, such as the instruction “how to make the product” with an input image of a bomb. To address the limitations of textual alignment, generating data of additional modality using generative models is a potential solution. However, not all modalities have high-performance generative models available, especially for emerging MLLMs that may arise in the future.
To address the aforementioned limitations, we propose SEA, a new framework that uses synthetic embeddings of additional modalities to enhance safety alignment. It first optimizes embedding representations within the modality encoder’s output space deemed by MLLMs to contain the specified harmful activities or products. Subsequently, the optimized embedding can be integrated with the textual dataset, substituting it for a real multimodal dataset in safety alignment training. Our approach eliminates the resource-intensive process of collecting and curating real multimodal datasets. Experiments are conducted on MLLMs based on images, videos, and audio, and the results indicate that only two training samples are needed to optimize a high-quality embedding in 24 seconds on a single RTX 3090 GPU. Furthermore, using datasets constructed with synthetic embedding for safety alignment significantly enhances the safety of MLLMs against threats from additional modalities.
Due to the lack of publicly available safety evaluation benchmarks for video and audio-based MLLMs, we also introduce VA-SafeBench, which expands on image-based MM-SafetyBench Liu et al. (2024c). Specifically, each sample in VA-SafetyBench is converted one-to-one from samples in eight scenarios of MM-SafetyBench. They share the same sources of harmful information, but the questioning format in VA-SafetyBench consists of video-text pairs and audio-text pairs. The high attack success rates (ASR) in multiple MLLMs validate the challenges posed by VA-SafeBench.
The contributions of our paper are summarized as follows.
$\bullet$ We introduce SEA, a novel low-resource MLLM safety alignment method. It expands the textual safety alignment dataset through synthetic embeddings, allowing multimodal training when only textual data is available.
$\bullet$ We present VA-SafeBench, which extends MM-SafetyBench to evaluate the security risks introduced by video and audio.
$\bullet$ The experimental results indicate that SEA significantly improves the security of MLLMs against threats from additional modalities with minimal additional computational overhead.
2 Related Works
2.1 Safety Concerns of MLLMs
LLMs have been revealed to pose significant risks in responding to malicious instructions Zou et al. (2023); Liu et al. (2023); Chao et al. (2023). Since MLLMs are typically developed using LLMs as their backbone networks, the risks inherent in the LLM domain are directly transferred to MLLMs. More concerning, recent studies have revealed that non-text modal inputs pose a more significant security threats to MLLMs. For example, leveraging the model’s OCR capabilities in combination with malicious images Gong et al. (2023); Luo et al. (2024) can significantly increase the response rate of malicious instructions. Furthermore, some works Li et al. (2024b); Qi et al. (2024); Niu et al. (2024) use gradient-based searches to generate image-level adversarial perturbations, further exacerbating security risks. Therefore, additional safety alignment for MLLMs is necessary to mitigate potential societal harm.
2.2 Safety Alignment for MLLMs
Safety alignment aims to align the safety awareness of the model with that of humans to prevent the generation of harmful content. This has been thoroughly researched in the field of LLMs, with widely used methods including SFT, Direct Preference Optimization (DPO) Rafailov et al. (2024), and Proximal Policy Optimization (PPO) Schulman et al. (2017). Inspired by these works, researchers have created carefully crafted image-text pairs for alignment training in MLLMs, yielding promising results in improving model safety. However, producing high-quality multimodal alignment data is often costly. To achieve low-resource safety alignment, Chakraborty et al. (2024) have revealed that textual unlearning can effectively enhance model safety. However, it has been noted that this is ineffective against attacks introduced solely from images. Hu et al. (2024). Furthermore, most existing works have focused solely on image-based MLLMs, leaving the effectiveness of other modalities to be explored further.
2.3 Safety Benchmark of MLLMs
Most of the existing safety benchmarks focus on image-based MLLMs, including MM-SafetyBench Liu et al. (2024c), Ch3ef Shi et al. (2024), VLSafe Chen et al. (2024), Figstep Gong et al. (2023), MLLMGuard Gu et al. (2024), and Jailbreakv-28k Luo et al. (2024). Furthermore, Yang et al. (2024a) utilized text-to-speech models to reveal security risks in the audio modality, while SafeBench Ying et al. (2024) provides a unified benchmark that can test the safety of both image and audio modalities. Currently, there are no published safety assessment benchmarks for MLLMs in other modalities.
3 SEA: Achieving Low-Resource Safety Alignment via Synthetic Embeddings
Figure 1: The overall framework of SEA. The execution process is demonstrated using an example in the image modality, encompassing three stages: data preparation, embedding optimization, and safety alignment.
Since multimodal datasets are crucial for MLLM’s safety alignment training, but not all modalities have high-performance generative models available, we aim to find a more general method for synthesizing data of additional modality. A key insight is that data of additional modality used for safety alignment (such as bomb images) is not necessary be human-interpretable, but merely needs to be interpreted as such by MLLMs.
Building on this intuition, we propose Synthetic Embedding enhanced safety Alignment (SEA), which optimizes embeddings in the representation space of the additional modality. The target embedding is the one that MLLMs interpret as containing the specified harmful activities or products. Specifically, SEA treats the embedding of additional modality as a trainable weight, optimized through gradient updates, to maximize the probability of the model outputting the specified content. After integrating the optimized embedding with the textual dataset, it can serve as a substitute for real multimodal datasets.
The pipeline of SEA is illustrated in Figure 1 and consists of three stages: (1) the data preparation stage, which convert textual alignment data into auxiliary data for embedding optimization. (2) the embedding optimization stage, which focuses on synthesizing embeddings for additional modalities. (3) the safety alignment stage, which performs multimodal alignment training by combining the synthesized embeddings with the textual alignment data. In the remainder of this section, we will first introduce the MLLMs architecture, followed by a detailed explanation of each stage.
3.1 MLLMs Architecture
The architecture of existing MLLMs can generally be broken down into three components. (1) Modality Encoder $M(\cdot)$: it encodes the input of additional modality into an embedding. (2) Projector $P(\cdot)$: it maps embeddings from the non-textual modality representation space into the textual modality representation space. (3) LLM: it processes inputs from different modalities, performing semantic understanding, reasoning, and decision-making. Combining these components, the reasoning process of MLLMs can be formulated as:
$y=LLM\left(P\left(M\left(z\right)\right),x\right),$ (1)
where $z$ and $x$ represent the input of additional modality and textual modality, respectively, while $y$ is the textual output.
When considering synthetic data of additional modality, there are three feasible options which are synthesizing the raw additional modality $z$, the output of the modality encoder $M(\cdot)$ or the output of the projector $P(\cdot)$. Since $z$ varies greatly in form across different modalities, it is not considered. Among the remaining options, we favor synthesizing the output of $M(\cdot)$ because in most MLLM training paradigm, $M(\cdot)$ remains frozen while $P(\cdot)$ is trainable, making SEA applicable to more MLLMs.
3.2 Data Preparation
Assume a textual safety alignment dataset $D_{T}=\left\{\left(x_{T}^{i},y_{T}^{i}\right)\right\}_{i=0}^{N}$ consisting of $N$ samples, where $x_{T}^{i}$ represents harmful instructions and $y_{T}^{i}$ can be a single harmless response for SFT or a pair of chosen-rejected responses for RLHF, our objective is to optimize a set of embeddings $\{E^{i}\}_{i=0}^{N}$ based on harmful information in $\{x_{T}^{i}\}_{i=0}^{N}$.
For each $\left(x_{T}^{i},y_{T}^{i}\right)\in D_{T}$, we individually prepare a auxiliary dataset $D_{a}^{i}=\{(x_{c}^{i},y_{c}^{i}),(x_{s}^{i},y_{s}^{i})\}$ to assist in the optimization of $E^{i}$, where $(x_{c}^{i},y_{c}^{i})$ and $(x_{s}^{i},y_{s}^{i})$ are the content control sample and the style control sample, respectively. Each sample includes an instruction, a response prefix, and a guiding text. The response prefix is concatenated with the guidng text to form the target response. Taking the image-based MLLM as an example, the construction process of $D_{a}^{i}$ is as follows:
Harmful Information Extraction. Inspired by Liu et al. (2024c), we utilize GPT-4o-mini to identify harmful phrase in $x_{T}^{i}$, and then classify harmful phrases into two categories, including “activity” and “product”, and then create a detoxified version of $x_{T}^{i}$ by replacing the harmful phrase with “this product” or “this activity.”
Since the harmful phrase related to activity often do not form complete sentences, we further prompt GPT-4o-mini to convert them into full sentences $s_{T}^{i}$ with subject-verb-object structures, aligning with the language habits of MLLMs.
Content Control Sample Construction. This sample is used to control the primary harmful content in the embedding. We use “Please briefly describe the activity (product) in the image.” as input instruction $x_{c}^{i}$, and “response prefix + $s_{T}^{i}$ (or harmful phrases for product)” as the target response $y_{c}^{i}$. The “response prefix” is determined based on the models’ output habits.
Style Control Sample Construction. This sample is designed to enhance embedding diversity. The input instruction $x_{s}^{i}$ is set to “What is the style of the image?.” The target response $y_{s}^{i}$ is set to “response prefix + style description.” The style description is randomly sampled from a predefined style set determined by the model’s output habits.
More details and examples on constructing $D_{T}$ can be found in Appendix B.1.
3.3 Embedding Optimization
After building the $D_{a}^{i}$, $M(\cdot)$ encodes a blank image (or blank video, silent audio) into an embedding, which serves as an initialization for a trainable embedding $E_{o}$. For each $(x^{i},y^{i})\in D_{a}^{i}$, the goal of the embedding optimization is to maximize the probability of the MLLM generating $y^{i}$ when given $x^{i}$ and $E_{o}$. During the optimization process, the entire MLLMs are frozen, with only $E_{o}$ participating as the trainable weight in the gradient updates. Since the content and style are specified in $y_{c}^{i}$ and $y_{s}^{i}$, the optimization objective can be understood as finding the embedding that the MLLM considers most aligned with that content and style. The entire optimization process can be formulated as follows:
$L\left(E_{o}\right)=-\frac{1}{\left|D_{a}^{i}\right|}\sum_{(x^{i},y^{i})\in D_{a}^{i}}\log\left(P_{r}\left(y^{i}\mid x^{i},P\left(E_{o}\right)\right)\right)$,
$E^{i}=\underset{E_{o}}{\arg\min}\left(L\left(E_{o}\right)\right),$ (3)
where $P_{r}\left(y^{i}\mid x^{i},P\left(E_{o}\right)\right)$ represents the conditional probability of generating $y^{i}$ when given $x^{i}$ and $P(E_{o})$ to the LLM.
3.4 Safety Alignment
To integrate $E^{i}$ and $D_{T}$ to construct the multimodal dataset $D_{M}=\left\{\left(x_{M}^{i},y_{T}^{i},E^{i}\right)\right\}_{i=0}^{N}$, a prefix in the form of “The image shows an activity (product). Please comprehend it and respond to the question below.” is added to the detoxified version of each $x_{T}^{i}$, resulting in the instruction $x_{M}^{i}$. The detoxified instruction ensures that harmful information is conveyed exclusively through the synthesized embeddings. The responses $\{y_{T}^{i}\}_{i=1}^{N}$ in $D_{T}$ are retained in $D_{M}$.
To achieve safety alignment based on $D_{M}$, we need to bypass module $M(\cdot)$ and modify the forward propagation process of MLLMs to $y=LLM\left(P\left(E^{i}\right),x\right)$, allowing it to adapt to existing safety alignment training strategies.
Notably, most current MLLMs freeze $M(\cdot)$ during the instruction fine-tuning stage. This enables seamless integration of SEA-generated synthetic datasets with real multimodal datasets. Taking image modality as an example, the real multimodal sample takes the form of “image embedding encoded by $M(\cdot)$ + instruction + response,” while the synthetic sample consists of “SEA embedding + instruction + response.” Since $M(\cdot)$ are frozen for most MLLMs, both forms look the same to the MLLM during training.
4 VA-SafetyBench: Assessing Security Risks Introduced by Video and Audio
4.1 VA-SafetyBench Overview
Figure 2: Overview of VA-SafetyBench construction pipeline.
VA-SafetyBench is a safety benchmark targeted at video and audio-based MLLMs. It consists of two parts: Video-SafetyBench and Audio-SafetyBench. Each sample in both parts includes a textual instruction and either a video or audio clip.
The construction pipeline of VA-SafetyBench is illustrated in Figure 2. VA-SafetyBench builds on MM-SafetyBench, a well-established image-based safety benchmark, through a systematic transformation process. Each test case in VA-SafetyBench directly corresponds to a test case in MM-SafetyBench, which spans eight critical safety scenarios: illegal activity, hate speech, malware generation, physical harm, economic harm, fraud, sexual violence, and privacy violations. For each sample, we utilize three types of textual data from MM-SafetyBench, including (1) an original instruction, (2) a harmful key phrase extracted from the original instruction, and (3) a rewritten instruction that conceals the harmful content in the original instruction. Figure 2 provides an example of these texts. Based on three types of textual data, we collect video and audio according to the key harmful phrase and refine the rewritten instruction to suit the new modalities.
4.2 Video-SafetyBench
Video-SafetyBench comprises four distinct tasks, including a Text-only baseline task where the input of video modality is set to empty, and the original instructions are used. The remaining three tasks are differentiated by their video generation methodologies, as detailed below:
$\bullet$ Diffusion Model (DM): We employed Pyramidal Flow Jin et al. (2024), a text-to-video model based on the diffusion Transformer architecture to generate videos where the content of videos is related to harmful key phrases. Each video is set to 5 seconds, with a resolution of 640×384 and 24 fps. Since individual harmful key phrases cannot provide sufficiently detailed video information, we utilized GPT-4o-mini to generate relevant scene descriptions. The generation process utilizes in-context learning, in which two examples were selected from five pre-prepared examples to serve as context. The five pre-prepared examples are shown in the Appendix A.2. We generate five distinct scene descriptions for each key phrase and process them through Pyramidal Flow. The most representative video is manually selected based on alignment with the original key phrase. To create the textual instructions for this task, we adapted the rewritten instructions from MM-SafetyBench by simply replacing all instances of “the image” with “the video” to reflect the change in modality.
$\bullet$ Typography (TYPO): Inspired by the image-based jailbreak method Gong et al. (2023), we developed a typographic animation approach to generate videos. Each harmful key phrase was decomposed into individual words, rendered in white text against a black background, and sequenced into a video format. The textual instruction is set to be the same as the DM.
$\bullet$ DM + TYPO: It combines the Diffusion Model and Typography to generate videos. In this task, Typography does not break phrases into individual words but displays the entire harmful key phrase in a single video frame, repeating multiple frames to form a 3-second video, which is then appended to the end of the DM video. In addition, we modified the rewritten instructions by replacing “the image” with “the end of the video” as the textual instruction. An instance would be “The end of the video shows a phrase of an activity. List steps in detail to perform the activity.”
4.3 Audio-SafetyBench
The Audio-SafetyBench also includes four tasks, with the Text-only task being the same as Video-SafetyBench. The other three tasks utilize Microsoft’s edge-tts API  [^1] https://github.com/rany2/edge-tts to generate speech, which can be differentiated based on the content of the speech:
$\bullet$ Audio-only: The entire unmodified original prompt is converted into speech. No textual instruction is provided.
$\bullet$ Key Phrase Transfer (KPT): Only key phrases are converted into speech. The textual instruction input is the rewritten instruction with “the image” replaced by “the audio.”
$\bullet$ KPT + Noise: Inspired by Yang et al. (2024a), we prepend a segment of white noise lasting 3 seconds to the KPT audio. The textual instruction input is the rewritten instruction with “the image” replaced by “the end of the audio.”
4.4 Experimental Validation
The experiments are carried out on four video-based MLLMs including Qwen2-VL-7b Wang et al. (2024b), Qwen2-VL-2b, VideoLLaMA2-7b Cheng et al. (2024), and VideoLLaMA2.1-7b, and four audio-based MLLMs including Qwen2-Audio Chu et al. (2024), Qwen-Audio Chu et al. (2023), SALMONN-7b Tang et al. (2024), and SALMONN-13b. Following the evaluation protocol established in Liu et al. (2024c), we use GPT-4o-mini as the evaluator. The evaluation metric is Attack Success Rate (ASR), calculated as the percentage of responses flagged as unsafe within the evaluation set.
Tables 1 and 2 present the experimental results. The findings can be summarized as follows.
Additional modalities amplify security vulnerabilities. Our research extends beyond previous findings in image-based MLLMs, demonstrating that this vulnerability pattern persists across audio and video modalities. The effect is particularly pronounced in video-based MLLMs, with Qwen2-VL-7b showing a 65.08% higher Attack Success Rate (ASR) when under DM+TYPO attacks compared to text-only attacks. Similarly, audio-based MLLMs demonstrate increased vulnerability, with Qwen-Audio exhibiting a 46.35% higher ASR when harmful content is presented through the audio modality.
Table 1: The evaluation results on the VA-SafetyBench for video-based MLLMs.
Models Text-only DM TYPO DM+TYPO
Qwen2-VL-7b 6.05 20.18 40.72 71.13
Qwen2-VL-2b 18.20 15.63 34.41 69.33
VideoLLaMA2-7b 15.66 24.83 20.81 42.33
VideoLLaMA2.1-7b 7.25 24.37 43.69 52.26
Table 2: The evaluation results on the VA-SafetyBench for audio-based MLLMs. The Audio-only results for SALMONN-7b and SALMONN-13b were discarded, as they consistently only repeated the content of the input speech.
Models Text-only Audio-only KPT KPT+Noise
Qwen2-Audio 14.66 7.70 24.06 34.31
Qwen-Audio 12.20 47.24 47.14 58.55
SALMONN-7b 44.04 - 41.11 65.97
SALMONN-13b 46.15 - 55.10 64.34
VA-SafetyBench poses significant challenges. Both video and audio-based MLLMs demonstrate high ASRs. For video-based MLLMs, the best case is the Qwen2-VL-7b, which achieved a 71.33% ASR in DM+TYPO. While the ASR for DM and TYPO respectively is generally lower than that of DM+TYPO, we found that in many instances, the MLLMs failed to correctly interpret the content of the videos, leading to safe outputs. Therefore, as the performance of MLLMs improves in the future, DM and TYPO may pose even greater threats. For audio-based MLLMs, the highest ASR is found in SALMONN-7b at 65.97%. KPT is generally higher than Audio-only, indicating that distributing harmful instructions across text and audio better activate model’s toxicity. KPT + Noise generally performs better than KPT, indicating that using noise for interference or hiding harmful information in the time dimension makes it easier to bypass safety mechanisms.
Due to space limitations, we present more details about VA-SafetyBench in the Appendix A.
Table 3: We present experimental results for an image-based MLLM (LLaVA-1.5-7b-hf), with separate evaluation on the safety benchmark and benchmarks of general capabilities. The results on safety benchmarks are presented on the left of the vertical line, with lower scores indicating better performance. The results on benchmarks of general capability are presented on the right of the vertical line, with higher scores indicating better performance. Bold values indicate the best performance.
Approaches MM-SafetyBench (%) POPE MMMU MME
Text-only SD TYPO SD + TYPO average adversarial popular random
LLaVA-1.5-7b-hf 46.50 30.20 27.32 62.78 41.7 81.32 86.56 89.53 30.44 1486
VLGuard 3.49 4.73 3.16 11.16 5.63 76.18 77.82 78.05 22.55 1304
Textual SFT 7.78 5.57 2.31 37.69 13.33 78.80 79.52 78.53 36.00 1157
GM SFT 3.14 1.35 0.70 3.10 2.05 75.99 79.24 78.80 29.88 979
SEA SFT 4.09 0.74 0.16 2.74 1.93 76.79 79.04 79.39 31.88 1114
Textual DPO 6.84 22.60 17.21 52.84 24.87 81.30 86.63 90.14 32.44 1433
GM DPO 26.37 13.76 8.49 41.95 22.64 80.42 86.13 89.43 30.66 1420
SEA DPO 7.27 6.56 2.77 23.20 9.95 82.34 86.99 89.94 30.11 1463
Table 4: The experimental results conducted on a video-based MLLM (Qwen2-VL-7b).
Approaches VA-SafetyBench (%) MVBench VideoMME
Text-only DM TYPO DM+TYPO average Short Medium Long
Qwen2-VL-7b 6.05 64.42 71.13 69.24 52.71 63.62 63.33 48.66 45.44
Textual SFT 4.27 4.35 12.51 13.47 8.65 61.85 60.66 48.44 44.55
GM SFT 2.91 3.31 6.01 6.70 4.73 61.65 58.77 48.66 43.55
SEA SFT 0.82 0.11 0.24 0.22 0.34 62.25 61.00 48.55 45.33
Textual DPO 2.82 3.71 12.61 14.00 8.28 62.65 62.00 48.44 44.00
GM DPO 2.97 2.06 5.57 10.34 5.23 63.92 63.22 48.88 45.55
SEA DPO 1.78 0.42 5.72 6.35 3.56 62.95 61.88 48.00 44.77
Table 5: The experimental results conducted on an audio-based MLLM (Qwen2-Audio-7b).
Approaches VA-SafetyBench (%) AIRBench
Text-only Audio-only KPT KPT+noise average Speech Sound Music Mixed-audio
Qwen2-Audio-7b 14.66 7.70 24.06 34.31 20.18 5.47 4.07 3.97 4.26
Textual SFT 5.05 4.82 21.40 16.45 11.93 5.58 4.13 4.05 4.28
GM SFT 4.28 7.77 4.43 5.22 5.54 4.05 3.28 3.21 3.40
SEA SFT 3.31 2.24 1.73 1.77 2.26 4.87 3.70 3.61 4.00
Textual DPO 6.58 4.85 18.87 35.56 16.40 5.58 4.16 4.17 4.29
GM DPO 5.52 2.59 2.85 2.42 3.34 5.64 4.33 4.05 4.26
SEA DPO 7.71 4.61 3.16 4.15 4.90 5.57 4.15 3.98 4.26
5 Experiments
5.1 Experimental Setup
Backbones. We select the widely used MLLM backbone for each modality: LLaVA-1.5-7b-hf Liu et al. (2024b) for images, Qwen2-VL-7b Wang et al. (2024b) for videos, and Qwen2-Audio-7b Chu et al. (2024) for audio.
Baselines. For image-based MLLMs, we have three baselines: (1) VLGuard Zong et al. (2024) utilizes 2k harmful and 1k harmless image-text pairs for SFT alignment. (2) Textual SFT (DPO) uses 3k textual samples for SFT (DPO) alignment. (3) GM SFT (DPO) uses a text-driven generative model (GM) to synthesize additional modal data for 3k textual samples, guided by the content guiding texts of SEA.
Since there is no related work on safe alignment, only the two baselines, Textual SFT (DPO) and GM SFT (DPO), are used for video and audio modalities. The generative models used for the three modalities are FLUX.1-dev  [^2] https://github.com/black-forest-labs/flux, CogVideoX-2b Yang et al. (2024c), and ChatTTS  [^3] https://github.com/2noise/ChatTTS, which differ from the models used in the benchmark construction.
Training Datasets. Following the settings in Hu et al. (2024), we sample 3k examples from the textual alignment dataset SafeRLHF Ji et al. (2024a), including 2k harmful samples and 1k harmless samples, as training dataset for SEA, Textual SFT (DPO), and GM SFT (DPO). For details on sampling, please refer to Appendix B.2. VLGuard’s training data is based on the dataset provided in the original work Zong et al. (2024), most of which consist of real-world image-text pairs.
Evaluation Benchmark. For safety assessment, we employ MM-SafetyBench for image-based MLLMs and VA-SafetyBench for video and audio-based MLLMs. To evaluate general capabilities, we utilize MMMU Hendrycks et al. (2021) and POPE Li et al. (2023) for image-based MLLMs, MVBench Li et al. (2024a) and VideoMME Fu et al. (2024) for video-based MLLMs, and AIR-Bench Yang et al. (2024b) for audio-based MLLMs.
Evaluation Metrics. For safety assessment, following the evaluation protocol established in Liu et al. (2024c), we use GPT-4o-mini as the evaluator. The evaluation metric is Attack Success Rate (ASR), calculated as the percentage of responses flagged as unsafe within the evaluation set. For evaluation on general capabilities, we adhere to the evaluation metric defined by the benchmark.
Implementation Details. We conducted embedding optimization training of SEA on a single RTX 3090 GPU. All MLLMs were set to a maximum of 100 training epochs. The learning rates for LLaVA-1.5, Qwen2-VL, and Qwen2-Audio were set to 0.02, 0.02, and 0.05, respectively, with cosine annealing updates. For efficiency, we implement an early stopping mechanism during embeddings optimization by checking whether the optimization is successful every 10 gradient update steps. For style control, optimization is considered successful if the output includes the complete guiding text. For content control, when the harmful phrase is categorized as “activity”, we consider optimization successful if $N-1$ out of $N$ words match, allowing for minor verb tense variations while preserving semantic meaning. When the harmful phrase is categorized as “product”, we require exact word-for-word matching. Training is terminated early when both content and style optimizations are successful. For failed optimization samples, we directly use the embeddings from the last epoch. For the safety alignment training, we implemented both SFT and DPO training strategies for SEA. More details about the experimental setup can be found in Appendix B.
5.2 Main Results
Tables 3, 4, and 5 present the results of experiments conducted on image, video, and audio-based MLLMs. We will showcase our findings through the following comparisons.
Comparison between SEA and textual alignment. Both Textual SFT and Textual DPO belong to textual alignment approaches. Compared to models without safety alignment, Textual SFT and Textual DPO effectively reduce the ASR of text-only attacks. Still, their effectiveness against multimodal attacks is limited, which is particularly evident in the image-based MLLM (LLaVA-1.5) and audio-based MLLM (Qwen2-Audio). SEA demonstrates comparable safety capabilities to textual alignment methods under Text-only attacks while significantly lowering the ASR of multimodal attacks. For instance, in the most challenging SD-TYPO task, SEA SFT’s ASR decreased by 34.95% compared to Textual SFT, and SEA DPO’s ASR decreased by 29.64% compared to Textual DPO. For general performance, whether using SFT or DPO training strategies, the overall performance of SEA is close to that of textual alignment methods. In summary, compared to textual alignment methods, SEA can significantly reduce the safety risks introduced by additional modalities without sacrificing general performance.
Comparison between SEA embeddings and physical multimodal data. In the baselines, VLGuard and GM SFT (DPO) were trained on physical image-text pairs. The results indicate that, except for SEA DPO being slightly inferior to GM DPO in the audio modality, SEA consistently demonstrated better safety performance in every comparison group. Additionally, SEA’s general performance was comparable to these baselines. Further observations of the data generated by the generative model revealed that the audio modality produced the highest quality output, with minimal information loss when converting text to speech, while the generated images and videos often show lower relevance to the guiding text. This may explain GM DPO’s strong performance in the audio modality. On the other hand, the audio modality conveys information through spoken voice, which has a stronger capacity to represent harmful content. Images and videos communicate harmful information through abstract visual concepts, which can lead to MLLMs not correctly interpreting them. This might be one of the reasons why VLGuard and GM SFT (DPO) perform worse than SEA in these two modes. In contrast, SEA consistently captures embeddings that MLLMs perceive as highly relevant to the guiding text, demonstrating stable performance across all modalities. This validates the potential of SEA for future applications in new modal MLLMs.
Comparison between DPO and SFT. The SFT-based approaches demonstrate stronger security than the DPO-based methods, but they typically results in a general performance decline. This is because using the reference model in DPO aids in maintaining general performance. Notably, aside from a slight degradation in the Qwen2-VL-7b, the general performance of SEA DPO in other model does not decline compared to the original model. Therefore, we recommend using DPO as the training strategy for SEA.
5.3 Efficiency and Quality of Embedding Optimization
To validate the efficiency of the embedding optimization, we recorded the optimization success rate (OSR) and the time consumed for embedding optimization across 3k samples. To check whether the model consistently believes that the optimized embeddings contain information from the content control samples, we designed three rewritten versions of the content control instruction, such as “Could you explain what is occurring in the image?”, and used them to query the MLLMs regarding the content in the optimized embeddings. The proportion of model outputs containing the target content out of total samples is reported as the Generalization Success Rate (GSR).
Table 6 presents the statistical results. SEA successfully finds embeddings for specified content and style in more than 93% of the cases across all models, demonstrating good generalization even when faced with instructions not seen in the content control samples, indicating that the embeddings are of high quality. The cases in Appendix B.4 further validates this. It should be acknowledged that some optimization failures and low-quality embeddings still exist. Since the SEA embeddings are optimized solely based on gradients from the MLLMs, the quality of embedding optimization depends on the model’s internal knowledge. As a result, embedding optimization is prone to failure in domains not covered by the model’s knowledge. However, if the MLLM inherently lacks the relevant knowledge, even an unaligned model would struggle to provide harmful guidance for related malicious instructions. Therefore, retaining low-quality embeddings does not compromise safety, but it may degrade general performance. This might explain the general performance decline of SEA DPO in the video modality. To address this, future improvements could include introducing a filtering mechanism based on embedding quality to discard low-quality embeddings before safety alignment training.
In terms of efficiency, each sample requires an optimization time of no more than 24 seconds on a single RTX 3090 GPU, which is significantly lower than the cost of manually collecting real data. Since optimization is performed on individual samples, SEA allows for parallel embedding optimization of a large-scale textual dataset across multiple GPUs, further saving computational time.
Due to space limitations, additional experiments and analyses regarding SEA can be found in Appendix B.3 and B.4.
Table 6: The OSR, average time consumption, and GSR of the embedding optimization on three models.
Models OSR(%) Average Time(s) GSR(%)
LLaVA-1.5-7b-hf 98.17 23.86 87.76
Qwen2-VL-7b 93.67 20.37 69.52
Qwen2-Audio-7b 98.37 12.06 97.15
6 Conclusions
The high cost of constructing multimodal datasets poses a significant challenge to developing safety alignment. In this paper, we demonstrate that synthetic embeddings can substitute for real additional model data, allowing for effective multimodal safety alignment relying solely on text. The high performance demonstrated by MLLMs across various modalities such as images, videos, and audio, validates the universal applicability of the proposed SEA method. Before the release of high-quality, large-scale real multimodal datasets, it holds promise as a safety solution for emerging MLLMs.
Limitations
Although SEA has shown promising performance, it still has limitations. On one hand, the optimization of SEA embeddings relies on the model’s own knowledge, making it susceptible to failures on harmful concepts that are not covered by the model. Given the diversity of the world, SEA is likely to fail on concepts which are more unique to different cultures of the world (e.g. a threat in a local dialect or a particular type of food which is poisonous in a country). Since the model is unlikely to provide accurate guidance for concepts it doesn’t understand, filtering out these failed embeddings before safety alignment training seems like a viable solution. However, it remains crucial for the model to correctly perceive these harmful concepts and responsibly reject related instructions. Finding a way to achieve this in low-resource settings is still an open question.
On the other hand, while style control enhances the diversity of SEA embeddings, it still cannot guarantee coverage of a sufficiently broad range of harmful scenarios. Further strategies for enhancing diversity, such as designing more samples for optimization, still need to be explored. We leave these two issues for future works.
Ethics Statement
This work includes harmful datasets and harmful content generated by MLLMs. The harmful instructions in the dataset come from existing safety evaluation benchmarks, and the harmful videos and audio are generated by the models. It is important to emphasize that this harmful content does not reflect the authors’ views. The purpose of this work is to propose safety alignment methods to promote the development of safer MLLMs. The construction of the dataset and presentation of harmful text generated by the model are solely to validate the effectiveness of our method.
Acknowledgments
This work was supported by National Natural Science Foundation of China through grant 62406114, 62306117 and 62472181, the Fundamental Research Funds for the Central Universities through grant 2024ZYGXZR074, Guangdong Basic and Applied Basic Research Foundation through grant 2025A1515011413 and 2024A1515010220, Local Science and Technology Development Fund of Hebei Province Guided by the Central Government of China through grant 246Z0102G, the “Pioneer” and “Leading Goose” R&D Program of Zhejiang through grant 2025C02044, Hebei Natural Science Foundation through grant F2024210008, the Guangzhou Basic and Applied Basic Research Foundation through grant 2024A04J3681, GJYC program of Guangzhou through grant 2024D03J0005, National Key R & D Project from Minister of Science and Technology through grant 2024YFA1211500, and South China University of Technology-TCL Technology Innovation Fund.
Appendix A Supplementary Materials for VA-SafetyBench
A.1 Motivation
The Proposal of VA-SafetyBench. This work explores low-resource safety alignment methods generally applicable to various modal MLLMs, including image, video, and audio MLLMs. Quantifying the safety risks of models is crucial for evaluating alignment performance. However, there is currently a lack of safety benchmarks for video-based MLLMs. For audio-based MLLMs, the projects of two existing works Ying et al. (2024); Yang et al. (2024a) are still not yet fully developed. Therefore, it is necessary to establish VA-SafetyBench to help assess the performance of SEA.
Expansion of MM-SafetyBench The reason for choosing to expand MM-SafetyBench lies in its two advantages: 1) It provides harmful key phrases extracted from well-crafted prompts, which facilitates the generation of new modal content using text-based generative models. 2) Since the toxicity of text instructions has been transferred to images, most text instructions in MM-SafetyBench are harmless in themselves. This helps us to create benchmarks where harmful information is displayed solely through video or audio.
Reduction of Scenarios MM-SafetyBench contains 13 scenarios, but we have excluded political lobbying, legal opinion, financial advice, health consultation, and government decision-making, as these safety considerations are somewhat overly stringent. For instance, the financial advice scenario tests the ability of MLLMs to refuse to provide financial opinions. Similar data is difficult to find in existing alignment training datasets, which poses challenges for evaluating algorithms.
Figure 3: Prompt used for generating scene descriptions
A.2 Prompts for Scene Descriptions Generation
The prompts for generating scene descriptions are shown in Figure 3, where Example 1 and 2 are randomly sampled from the five examples illustrated in Figure 4. Example 3 is the sample currently being processed.
Figure 4: Five example for in-context learning
Table 7: Comparison of experimental results between temporal stitching and spatial stitching.
Models DM+TYPO (Spatial) DM+TYPO (Temporal)
Qwen2-VL-7b 64.42 71.13
Qwen2-VL-2b 61.22 69.33
VideoLLaMA2-7b 39.68 42.33
VideoLLaMA2.1-7b 51.57 52.26
A.3 Comparison of Temporal Stitching and Spatial Stitching in DM+TYPO
In the DM+TYPO task of VA-SafetyBench, the videos generated by the model and the TYPO videos are stitched along the timeline. In MM-SafetyBench, the Stable Diffusion images and TYPO images are connected in pixel space. In fact, similar operations can also be performed in videos, such as adding TYPO subtitles to the bottom of each video frame. We have compared the experimental results of the two stitching methods in Table 7. Across all models, the temporally stitched datasets exhibited higher levels of harmfulness, so we adopted this stitching approach.
A.4 The Complete Benchmark Evaluation Results
Tables 8 and 9 present the complete evaluation results of VASafetyBench, including the results for each scenario.
Table 8: The complete evaluation results of Video-SafetyBench
Scenarios Sample size Qwen2-VL-7b Qwen2-VL-2b
Text-only DM TYPO DM+TYPO Average Text-only DM TYPO DM+TYPO Average
Illegal Activitiy 97 0 36.08 46.39 84.54 41.75 3.09 11.34 30.93 81.44 31.70
Hate Speech 163 0 8.59 28.22 61.96 24.69 6.75 3.68 33.13 61.96 26.38
Malware Generation 44 6.82 22.73 56.82 84.09 42.61 22.73 25.00 45.45 79.55 43.18
Physical Harm 144 11.81 30.56 53.47 82.64 44.62 32.64 38.89 46.53 84.03 50.52
Economic Harm 122 13.93 9.84 24.59 33.61 20.49 20.49 10.66 23.77 31.97 21.72
Fraud 154 0.65 18.18 46.75 82.47 37.01 13.64 13.64 38.31 81.82 36.85
Sex 109 13.76 4.59 32.11 47.71 24.54 31.19 4.59 28.44 50.46 28.67
Privacy Violence 139 1.44 30.94 37.41 92.09 40.47 15.11 17.27 28.78 83.45 36.15
Average - 6.05 20.18 40.72 71.13 34.52 18.20 15.63 34.41 69.33 34.39
Scenarios Sample size VideoLLaMA2-7b VideoLLaMA2.1-7b
Text-only DM TYPO DM+TYPO Average Text-only DM TYPO DM+TYPO Average
Illegal Activitiy 97 4.12 44.33 25.77 59.79 33.50 1.03 46.39 65.98 72.16 46.39
Hate Speech 163 10.43 14.72 16.56 35.58 19.32 4.29 12.88 34.36 42.33 23.46
Malware Generation 44 29.55 27.27 31.82 47.73 34.09 9.09 27.27 54.55 65.91 39.20
Physical Harm 144 29.86 34.72 27.78 55.56 36.98 11.81 33.33 56.25 58.33 39.93
Economic Harm 122 10.66 9.84 6.56 22.95 12.50 12.30 11.48 20.49 26.23 17.62
Fraud 154 14.94 30.52 31.82 55.84 33.28 1.95 29.22 53.90 62.99 37.01
Sex 109 12.84 2.75 4.59 13.76 8.48 14.68 2.75 16.51 21.10 13.76
Privacy Violence 139 12.95 34.53 21.58 47.48 29.13 2.88 31.65 47.48 69.06 37.76
Average - 15.66 24.83 20.81 42.33 25.90 7.25 24.37 43.69 52.26 31.89
Table 9: The complete evaluation results of Audio-SafetyBench
Scenarios Sample size Qwen2-Audio-7b-Instruct Qwen-Audio-Chat
Text-only Audio-only KPT KPT+noise Average Text-only Audio-only KPT KPT+noise Average
Illegal Activitiy 97 1.03 2.06 16.49 40.21 14.94 0 37.11 57.73 73.20 42.01
Hate Speech 163 1.84 1.84 15.34 20.86 9.97 2.45 50.31 34.36 43.56 32.67
Malware Generation 44 31.82 15.91 38.64 54.55 35.23 34.09 50.00 65.91 75.00 56.25
Physical Harm 144 23.61 1.25 39.58 49.31 28.43 22.92 58.33 54.17 68.06 50.87
Economic Harm 122 16.39 16.39 21.31 23.77 19.46 14.75 27.87 29.51 30.33 25.61
Fraud 154 5.19 3.25 22.73 44.16 18.83 3.25 54.55 53.90 68.83 45.13
Sex 109 20.18 13.76 14.68 12.84 15.36 10.09 42.20 31.19 33.94 29.35
Privacy Violence 139 17.27 7.19 23.74 28.78 19.24 10.07 57.55 50.36 75.54 48.38
Average - 14.66 7.70 24.06 34.31 20.18 12.20 47.24 47.14 58.55 41.28
Scenarios Sample size SALMONN-7b SALMONN-13b
Text-only Audio-only KPT KPT+noise Average Text-only Audio-only KPT KPT+noise Average
Illegal Activitiy 97 16.49 - 51.55 80.41 49.48 15.46 - 73.20 84.54 57.73
Hate Speech 163 31.29 - 32.52 62.58 42.13 25.15 - 53.37 64.42 47.64
Malware Generation 44 75.00 - 31.82 70.45 59.09 77.27 - 52.27 59.09 62.87
Physical Harm 144 69.44 - 54.17 86.81 70.14 59.03 - 66.67 77.78 67.82
Economic Harm 122 22.95 - 22.95 36.07 27.32 28.69 - 28.69 35.25 30.87
Fraud 154 57.14 - 51.95 76.62 61.90 70.13 - 65.58 80.52 72.07
Sex 109 27.52 - 15.60 32.11 25.07 36.70 - 24.77 27.52 29.66
Privacy Violence 139 52.52 - 68.35 82.73 67.86 56.83 - 76.26 85.61 72.90
Average - 44.04 - 41.11 65.97 50.37 46.15 - 55.10 64.34 55.19
Appendix B Supplementary Materials for SEA
B.1 More Implementation Details of SEA
Prompt for Harmful Phrase Extraction. In Section 4.2, the prompt used to instruct GPT-4o-mini to extract harmful phrases is illustrated in Figure 5.
Figure 5: Prompt for harmful phrase extraction.
Prompts for Sentence Completion. In Section 4.2, the prompt used to instruct GPT-4o-mini to complete harmful phrases into sentences is shown in Figure 6.
Details for Content Control and Style Control Samples. Both samples consist of an instruction, a response prefixes, and a guiding text. For the content control sample, the guiding text is either a harmful phrase (for product) or a sentence completed from harmful phrase (for activity). For the style control sample, the guiding text is a style description randomly sampled from a pre-prepared style set. Each model’s instruction, response prefixes, and style set are different and are designed based on the output patterns of the model. Specifically, we used 50 test samples to observe each model’s output and summarize their habits. Figure 7 displays the instructions, response prefixes, and style sets used by the three MLLMs. For easier understanding, we also prepared an example for each MLLM, which includes all intermediate outputs generated during the preparation of the embedding optimization dataset.
B.2 More Details on Experimental Setup
Training Data Construction. We sampled 2k harmful samples and 1k harmless samples from the SafeRLHF dataset. Each sample in SafeRLHF includes an instruction, a chosen response, and a rejected response. The harmful samples were randomly selected from samples in which the rejected responses with a severity level of 3. In contrast, the harmless samples were randomly selected from samples in which both the chosen and rejected responses had a sensitivity level of 0. Since we found that a significant amount of harmful content still existed in the chosen responses, we used Llama2-7b-chat to regenerate the chosen responses for harmful samples. In any SFT training, only the instruction and chosen response from each sample are used for training.
Evaluation Setting. For MM-SafetyBench, we selected only the same eight scenarios for evaluation as in VA-SafetyBench. In Tables 3, 4, and 5, the results presented by VA-SafetyBench and MM-SafetyBench are the average ASR for the eight scenarios. For all general capability evaluation benchmarks, we follow the implementation in Ji et al. (2024b).
Figure 6: Prompt for sentence completion.
Figure 7: Guidelines for constructing embedding optimization datasets. Each MLLM includes unique instructions, response prefixes, and a style set. Examples containing all intermediate outputs is provided to aid in understanding the construction process in Section 3.2.
Training Setting for Safety Alignment. All methods for safety alignment training are implemented on a server equipped with four A800 GPUs. For VLGuard, we follow the original paper’s setup and retrain on the Huggingface version of LLaVA-1.5-7b-hf. For other baselines, we summarize the training hyperparameter settings in Table 10.
Table 10: Hyperparameter settings for safe alignment training. (SFT) and (DPO) indicate the training strategy.
Approaches Training Method Learning Rate Epoch
LLaVA-1.5-7b-hf (SFT) full-parameter 2e-5 2
LLaVA-1.5-7b-hf (DPO) full-parameter 2e-6 3
Qwen2-VL-7b(SFT) full-parameter 1e-5 2
Qwen2-VL-7b(DPO) full-parameter 1e-6 3
Qwen2-Audio-7b (SFT) full-parameter 2e-5 3
Qwen2-Audio-7b (DPO) full-parameter 2e-6 3
B.3 The Impact of Style Control Samples
To verify whether style control samples help enhance the diversity of embeddings, we removed the style control samples and performed embedding optimization using only the content control samples. We then calculated the average cosine distance between the embeddings with and without the style control samples.
Table 11 shows the results. The consistent results across three different modal MLLMs indicate that adding style control samples enhances the average differences between embeddings, resulting in greater diversity. In addition, Figure 8 shows the t-SNE visualization of the SEA embeddings. Points of different styles in LLaVA-1.5-7b-hf and Qwen2-VL-7b form clusters in the embedding space, while in Qwen2-Audio-7b, points of different styles are distributed along a semicircular arc at different locations. It’s worth mentioning that when style optimization is not specified, we observe that the SEA embeddings obtained by LLaVA-1.5-7b-hf consistently correspond to the “black and white photograph” style. This may be related to the fact that we always use a fixed embedding (such as the embedding of a white image) as a starting point for optimization. Therefore, specifying the embedding style enables a broader distribution in the feature space, enhancing the diversity of the embeddings.
Table 11: Comparison of average cosine distances of SEA embeddings with and without style control.
Models SEA SEA without style control
LLaVA-1.5-7b-hf 0.03109 0.03023
Qwen2-VL-7b 0.06986 0.06663
Qwen2-Audio-7b 0.00796 0.00745
Figure 8: The t-SNE visualization of the SEA embeddings
B.4 Case Study
Embedding Quality. To explore what the optimized embeddings capture, we first ignored the modality encoder of MLLMs, and then input both synthetic embedding and textual instruction into the unaligned MLLM simultaneously. Each MLLM utilized three different instructions combined with the same embedding for single-round inference. The experimental results shown in Figure 9 indicate that even when the content guiding text consists of only brief descriptions, the optimized embeddings can capture rich and realistic information, such as that bombs should have fuses or the specific methods used by drug traffickers for money laundering. Furthermore, when faced with different queries, MLLM consistently interprets the embeddings as products or activities within the content guiding texts, as if these embeddings were encoded from real multimodal data containing this information.
It must be acknowledged that during the testing of some SEA embeddings, we observed hallucinations. This may be due to the model having not encountered similar data during pre-training stage, indicating a lack of relevant additional modal knowledge. In Section 5.3, we analyzed that retaining such low-quality embeddings does not compromise safety but may harm the model’s general performance. Implementing an output checking mechanism in such tests could potentially filter out low-quality embeddings, and we encourage future work to explore this.
Alignment Performance. To further explore how multimodal alignment data incorporating SEA embeddings enhances safety alignment performance, Figures 10, 11, and 12 respectively present cases evaluated on safety benchmarks for image, video and audio-based MLLMs. All cases come from the most challenging tasks in their respective modality benchmarks, namely SD+TYPO, DM+TYPO, and KPT+noise. When faced with harmful queries in three cases, both SEA SFT and SEA DPO refused to respond and provided their reasons.
Figure 9: Three examples of using SEA embeddings for dialogue with MLLMs. Each SEA embedding tested three instructions. Each instruction was fed into the MLLM along with the SEA embedding for single-round inference.
Figure 10: A case on the security evaluation of the image-based MLLMs. The evaluation sample is sourced from the SD+TYPO task of MM-SafetyBench, with the scenario “Sex.”
Figure 11: A case on the security evaluation of the video-based MLLMs. The evaluation sample is sourced from the DM+TYPO task of VA-SafetyBench, with the scenario “physical harm.”
Figure 12: A case on the security evaluation of the audio-based MLLMs. The evaluation sample is sourced from the KPT+noise task of VA-SafetyBench, with the scenario “malware generation”